OpenStudy (anonymous):

Find the number of sides of a convex polygon if the measures of its interior angles have a sum of 12,600 degrees

OpenStudy (anonymous):

sum of external angles = 360 so each ext angle = 360/n and interior angle = 180 - (360/n) where n = no. of sides so n ( 180 - 360/n) = 12600 solve this for n

OpenStudy (anonymous):

solving 180n - 360 = 12600 180n = 12960 n = 12960/180 = 72 sides

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

How did you know what formula to use?

OpenStudy (anonymous):

i started from the fact that sum of external angles = 360 degrees perhaps a drawing will help

OpenStudy (anonymous):



OpenStudy (anonymous):

the 6 exterior angles e add up to 360 degrees so each ext angles = 60 degrees an interior angle i is 180 - e = 120 so total angles in the figure = 6 * 120 = 720

OpenStudy (anonymous):

Okay, now I understand. Thank you

OpenStudy (anonymous):

there is a formula for the total number od degrees is polygon of no. sides = n it is total degrees = 180(n - 2) so here 12600 = 180( n - 2) solve for n

OpenStudy (anonymous):

12600 = 180n - 360 180n = -12600 - 360 180n = 12960 n = 72

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i forget that formula for some reason ! - its easier using that

OpenStudy (anonymous):